Solve the method of separation of variables du/dx+u=du/dt if u=4e–3x when t= 0
1
Expert's answer
2021-08-12T08:08:20-0400

Solution

Let u(x,t) = X(x)T(t)

Substituting u back into equation and dividing by X(x)T(t)

"\frac{X'}{X}+1=\frac{T'}{T}"

Since the right hand side depends only on x and the left hand side only on t, both sides are equal to some constant value −C. Thus,

"\frac{X'}{X}+1=-C,\ \ \ \ \frac{T'}{T}=-C"

From this ODE: X(x) = Ae-(C+1)x, T(t) = Be-Ct  , u(x,t) = De-(C+1)x-Ct (A,B,C,D  are arbitrary constants)

From initial condition u(x,0)=4e–3x  =>  De-(C+1) = 4e–3x  =>  D=4, C=2

Therefore the solution is u(x,t) = 4e-3x-2t